OpenStudy (anonymous):

I have a trapezoid. WX is parallel to YZ and points A and B are midpoints of their respective legs... If WX = 5x + 3, ZY= 13x - 1, and AB= 6x+7, find x

OpenStudy (anonymous):

I keep getting a fraction

OpenStudy (amistre64):

wx = yz 5x+3 = 13x -1 4 = 8x x = 2

OpenStudy (amistre64):

trapeziod doesnt imply equality of sides ....missed that one lol

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

fractions are good as answers you know :)

OpenStudy (anonymous):

midpoints of two sides is one-half the length of the third...

OpenStudy (anonymous):

Fractions complicate things! LOL

OpenStudy (amistre64):

:) a picture would be good

OpenStudy (anonymous):

it looks like a trapezoid sitting on top of another trapezoid.. getting smaller to the top (shorter lines)

OpenStudy (anonymous):

I think: [( 5x + 3)+(13-1)]/2=6x+7?

OpenStudy (anonymous):

sorry 23x not 13

OpenStudy (anonymous):

[( 5x + 3)+(13x-1)]/2=6x+7?

OpenStudy (anonymous):

ok I'll see what I get...

OpenStudy (anonymous):

x = 7/3 it cant be a fraction can it???

OpenStudy (anonymous):

x=2?

OpenStudy (anonymous):

I redid it... yes x = 2 i divided the 2 instead of multiplied it to cancel it out!! Awesome thank you both!! You both came up with the same answer!!!

OpenStudy (anonymous):

If I deserve medal then give me one.

OpenStudy (anonymous):

You GOT IT!!!!

OpenStudy (anonymous):

thank

OpenStudy (anonymous):

Thanks you too!!

OpenStudy (anonymous):

(18x+2)/2=6x+7 9x+1=6x+7 9x -6x=7-1 3x = 6 x =6/3 x =2

OpenStudy (anonymous):

Yep Got it!!! Great job!